# Contrast-Aware Calibration for Fine-Tuned CLIP: Leveraging Image-Text Alignment

### v1

@@ 1 Introduction
Vision-language models, such as CLIP (Radford et al., 2021), pre-trained on vast web-scale text-image datasets, have demonstrated impressive zero-shot capabilities and image-text alignment in various downstream image classification tasks (Deng et al., 2009; Helber et al., 2019). Various prompt learning methods have been proposed to enhance VLM performance on specific tasks using limited labeled data (Zhou et al., 2022b; Khattak et al., 2023a, b). Given CLIP’s strong zero-shot adaptability, the open-vocabulary setting has become a standard for evaluating the performance of fine-tuned VLMs, where prompts are trained on a subset of classes and evaluated on both train and unseen classes (Lee et al., 2023; Tan et al., 2024).
@@ 2 Related Works
Confidence Calibration. Confidence calibration has been widely studied to ensure that the confidence levels output by models accurately reflect their true performance. A common approach to achieving this is by applying calibration techniques after the model has been trained. These calibration methods can typically be categorized into two main paradigms: scaling-based methods (Guo et al., 2017; Tomani et al., 2022; Yu et al., 2022; Xiong et al., 2023) and bin-based methods (Zadrozny & Elkan, 2001; Zhang et al., 2020; Zadrozny & Elkan, 2002; Roelofs et al., 2022). For scaling-based calibration, a representative method is temperature scaling (Tomani et al., 2022), which learns a single temperature value for rescaling the logits. ATS (Joy et al., 2023) modifies the confidence predicted by per-data-point adaptive temperature. With the wide use of VLMs, several works have explored calibration in VLMs (LeVine et al., 2023; Oh et al., 2023) recently. Distance-Aware Calibration (Wang et al., 2024) estimates the scaling weights for unseen class logits based on textual features, focusing solely on changes in the textual modality. As a result, its ability to handle confidence calibration is limited when encountering fine-tuned VLMs with well-aligned image-text features. In this paper, we propose a novel confidence calibration method, leveraging the image-text alignment of CLIP.
@@ 6.1 Experimental Setup
Compared Methods. We mainly focus on benchmarking against the other 5 current representative prompt learning methods: CoCoOp (Zhou et al., 2022a), KgCoOp (Yao et al., 2023), MaPLe (Khattak et al., 2023a), ProGrad (Zhu et al., 2023), and PromptSRC (Khattak et al., 2023b). Since models like CoOp only consider the text modality and exhibit low accuracy, making their calibration significance minimal, we focused on testing and calibrating its optimized version, CoCoOp and KgCoOp. For train classes calibration, we select three representative calibration methods: Histogram Binning (HB) (Zadrozny & Elkan, 2001), Isotonic Regression (IR) (Zadrozny & Elkan, 2002), and Multi-Isotonic Regression (MIR) (Roelofs et al., 2022). For unseen classes calibration, we compare the SOTA method designed for CLIP, DAC (Wang et al., 2024). Datasets and Implementation details used in the experiments are shown in Appendix A.
Implementation details. In the main experiments of this paper, we set $k=15$ and $\alpha=1.10$ as the default parameters and select $\lambda_{1}=0.9$ and $\lambda_{2}=1.0$ as the threshold values. Additional details about pre-trained models, hyperparameters, and implementation specifics are provided in Appendix A.
@@ Appendix A Experimental Setting
Dataset. Following the CoCoOp framework (Zhou et al., 2022a), we conducted evaluations of our proposed DAC along with comparison calibration methods on various image classification tasks. These tasks included general object recognition using ImageNet (Deng et al., 2009) and Caltech101 (Fei-Fei et al., 2004) datasets, fine-grained object recognition involving datasets such as Oxford Pets  (Parkhi et al., 2012), Food-101 (Bossard et al., 2014), StanfordCars (Krause et al., 2013), Oxford Flowers102 (Nilsback & Zisserman, 2008), and FGVCAircraft (Maji et al., 2023). Additionally, we performed a remote sensing recognition task using the EuroSAT  (Helber et al., 2019) dataset, a texture recognition task using the DTD (Cimpoi et al., 2014) dataset, an action recognition task using UCF101 (Soomro et al., 2012) dataset and a large-scale scene understanding task using SUN397 (Xiao et al., 2010) dataset.
Implementation details. We use CLIP (ViT-B/16) as the pre-trained VLM throughout our experiments and report results averaged over 3 runs. We fine-tune the model with 16 samples per class in a few-shot setting (Zhou et al., 2022b). Following the official implementation, We list the general hyperparameters in Table 7. Here, we briefly introduce the corresponding exclusive hyperparameters of each VLM tuning method. All the methods are adopted from their official implementation. For CoCoOp, they do not contain other hyperparameters. For ProDA, we set $\lambda=0.1$ For KgCoOp, we set $\lambda=8.0$. For MaPLe, we set prompt depth $J$ to 0 and the language and vision prompt lengths to 2. For ProGrad, we set $\lambda=0.8$. For PromptSRC, we set deep prompting with V = T = 4. $\lambda_{1}=10$ and $\lambda_{2}=25$ are used in weight loss. For textual diversity, we use a total of $N=60$ standard prompt templates. We reproduce their methods in our framework based on the standard codebase.
@@ Appendix B The Impact of Different backbones
Follow DAC (Wang et al., 2024) experimental settings, we evaluated the performance of DAC and CAC on fine-tuned VLMs with different architectures using the OxfordFlowers (Nilsback & Zisserman, 2008) dataset in this section. Unlike accuracy, which tends to increase as the number of shots increases (Zhou et al., 2022a), calibration errors are prevalent across diverse few-shot settings. As shown in Table 8, CAC demonstrates stronger cross-model generalization capabilities, achieving effective calibration across all architectures.
Table 8: Comparison results of ECE (%) using different visual backbones on OxfordFlowers dataset. The smaller values are better.
@@ Appendix C Detailed Experimental Results
In this section, we illustrate the experimental results in detail.
@@ C.1 The calibration performance of fine-tuned CLIP
To showcase the miscalibration in fine-tuned VLMs, we fine-tune the pre-trained CLIP with 5 tuning methods on 11 downstream datasets and use ECE as the calibration metric. We report the detailed calibration results in this section. As is shown in Table 9(b), we can observe that distillation-based tuning (KgCoOp, ProGrad and PromptSRC) restrains the model’s confidence in train classes. For the calibration of the unseen class, distillation-based tuning utilizes zero-shot CLIP as the teacher model to improve the generalization ability, which generates relatively reliable predictions in unseen classes as expected. More importantly, we observe that fined-tuned VLMs tend to be underconfident in train classes while prone to overconfident in unseen classes. This phenomenon motivates us to investigate the calibration of fine-tuned VLMs further.
Table 9: Expected Calibration Error (ECE) on diverse downstream datasets using various tuning methods for CLIP-ViT-B/16. The calibration performance is averaged across three variants. Caltech, OP, SCars, Flowers, Food, FGVC, SUN, and UCF represent Caltech101, OxfordPets, StanfordCars, Flower102, Food101, FGVCAircraft, SUN397, and UCF101.
(a) Base
@@ C.2 Calibrated performance for fine-tuned CLIP
We utilize commonly used post-hoc confidence calibration for fine-tuned CLIP and expect to calibrate them well on both train and unseen classes. We can observe that established calibration can remedy miscalibration in train classes. However, such efficacy can not transfer to unseen classes.
@@ C.2.1 TRAIN CLASS
Post-hoc calibration can remedy miscalibration in train classes. In the latest works (Zadrozny & Elkan, 2001, 2002), scaling-based methods like TS have been successfully adopted in VLM calibration, and their findings reveal that this calibration can handle miscalibration well in the close-world setting. As a complement, we observe that bin-based (e.g., MIR) calibration can notably reduce miscalibration in train classes. Post-hoc calibration on train classes can not be transferred to unseen classes. For instance, scaling-based methods like TS scale the logits by a single scalar temperature value. If fine-tuned VLMs show underconfidence in train classes, TS sharpens the logit scale before the softmax function and makes the prediction more confident. As a cost, the confidence distribution of fine-tuned VLMs become sharper and causes these models to display exaggerated overconfidence in unseen classes. for bin-based methods, they need the probabilities from train classes as the input, which are incompatible with zero-shot prediction.
@@ C.2.2 UNSEEN CLASS
Table 10: Expected Calibration Error (ECE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViTB/16, with calibration performance averaged across three variants.
@@
Table 11: Expected Calibration Error (ECE) on train classes in diverse downstream datasets using various tuning methods for CLIP-ViTB/16, with calibration performance averaged across three variants.
@@ C.3 Accuracy & Confidence of fine-tuned CLIP
In this section, we analyze the miscalibration from the classification performance. Since open-vocabulary classes are not explicitly seen during the tuning, some might believe that this miscalibration arises from rapid accuracy decreasing and leads to a larger Expected Calibration Error (ECE). We report the average performance on 11 datasets in Table 12(b). Although the classification performance on unseen classes may be comparable or lower, the average predictive confidence of fine-tuned VLMs is still relatively higher than zero-shot CLIP, which results in a larger ECE. The results are the same as for DAC, although fine-tuning improves the accuracy of VLM in downstream tasks, the corresponding ECE also increases.
@@ C.4 Detailed results of main experiment
In this section, We present the detailed results on the calibration of unseen classes to verify that CAC can open-vocabulary calibration in existing prompt tuning. For comprehensive evaluation, 4 standard metrics are used in our evaluation of open-vocabulary confidence calibration: Expected Calibration Error (ECE) (Guo et al., 2017), Maximum Calibration Error (MCE) (Guo et al., 2017), Adaptive Calibration Error (ACE) (Nixon et al., 2019) and Proximity Informed Expected Calibration Error (Xiong et al., 2023). ALl of calibration error is given by ×$\times 10^{-2}$.
Table 14: Expected Calibration Error (ECE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViTB/16, with calibration performance averaged across three variants.
@@
Table 15: Adaptive Calibration Error (ACE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViTB/16, with calibration performance averaged across three variants.
@@
Table 16: Maximum Calibration Error (MCE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViTB/16, with calibration performance averaged across three variants.
@@
Table 17: Proximity-Informed Expected Calibration Error (PIECE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViTB/16, with calibration performance averaged across three variants.

### v3

@@ 1 Introduction
Vision-language models, such as CLIP (Radford et al., 2021), pre-trained on vast web-scale text-image datasets, have demonstrated impressive zero-shot capabilities and image-text alignment in various downstream image classification tasks (Deng et al., 2009; Helber et al., 2019). Various prompt learning methods have been proposed to enhance VLM performance on specific tasks with a small amount of labeled data (Zhou et al., 2022b; Khattak et al., 2023a, b). Given CLIP’s strong zero-shot adaptability, the open-vocabulary setting has become a standard for evaluating the performance of fine-tuned VLMs, where prompts are trained on a subset of classes and evaluated on both train and unseen classes (Lee et al., 2023; Tan et al., 2024).
@@ 2 Related Works
Confidence Calibration. Confidence calibration aims to align the confidence scores predicted by models with their actual performance. A common strategy for achieving this is to apply calibration techniques after model training. These techniques can be broadly divided into two categories: scaling-based methods (Guo et al., 2017; Tomani et al., 2022; Yu et al., 2022; Xiong et al., 2023) and bin-based methods (Zadrozny & Elkan, 2001; Zhang et al., 2020; Zadrozny & Elkan, 2002; Roelofs et al., 2022). Among the scaling-based approaches, temperature scaling (Tomani et al., 2022) is widely used, where a single temperature parameter is learned to adjust the logits. ATS (Joy et al., 2023), adapts the temperature for each data point individually. With the growing popularity of VLMs, recent studies have also examined the calibration of these models (LeVine et al., 2023; Oh et al., 2023). Distance-Aware Calibration (Wang et al., 2024) estimates the scaling weights for unseen class logits based on textual features, focusing solely on changes in the textual modality. As a result, its ability to handle confidence calibration is limited when encountering fine-tuned VLMs with well-aligned image-text features. In this paper, we propose a novel confidence calibration method, leveraging the image-text alignment of CLIP.
@@ 6.1 Experimental Setup
Compared Methods. We mainly focus on benchmarking against the other 5 current representative prompt learning methods: CoCoOp (Zhou et al., 2022a), KgCoOp (Yao et al., 2023), MaPLe (Khattak et al., 2023a), ProGrad (Zhu et al., 2023), and PromptSRC (Khattak et al., 2023b). Since models like CoOp only consider the text modality and exhibit low accuracy, making their calibration significance minimal, we focused on testing and calibrating its optimized version, CoCoOp and KgCoOp. For train classes calibration, we select three representative calibration methods: Histogram Binning (HB) (Zadrozny & Elkan, 2001), Isotonic Regression (IR) (Zadrozny & Elkan, 2002), and Multi-Isotonic Regression (MIR) (Roelofs et al., 2022). For unseen classes calibration, we compare the SOTA method designed for CLIP, DAC (Wang et al., 2024). Datasets used in the experiments are shown in Appendix A.
Implementation details. For the main results, we use CLIP (ViT-B/16) as the pre-trained VLM throughout our experiments and report results averaged over 3 runs. For the compared methods, we use their official implementations. In the main experiments of this paper, we set $k=15$ and $\alpha=1.10$ as the default parameters for our Contrast-Aware Weight. We selected $\lambda_{1}=0.9$ and $\lambda_{2}=1.0$ as the threshold values for the piecewise function. The rationality of these parameter choices will be validated through ablation experiments. Additional details about pre-trained models, hyperparameters, and implementation specifics are provided in Appendix A.
@@ Appendix A Experimental Setting
Dataset. Following the CoCoOp framework (Zhou et al., 2022a), we conducted evaluations of our proposed CAC along with comparison calibration methods on various image classification tasks. These tasks included general object recognition using ImageNet (Deng et al., 2009) and Caltech101 (Fei-Fei et al., 2004) datasets, fine-grained object recognition involving datasets such as Oxford Pets  (Parkhi et al., 2012), Food-101 (Bossard et al., 2014), Stanford Cars (Krause et al., 2013), Oxford Flowers102 (Nilsback & Zisserman, 2008), and FGVCAircraft (Maji et al., 2023). Additionally, we performed a remote sensing recognition task using the EuroSAT  (Helber et al., 2019) dataset, a texture recognition task using the DTD (Cimpoi et al., 2014) dataset, an action recognition task using UCF101 (Soomro et al., 2012) dataset and a large-scale scene understanding task using SUN397 (Xiao et al., 2010) dataset.
Implementation details. In our experiments, we utilize CLIP (ViT-B/16) as the pre-trained VLM and report results averaged over 3 runs. The model is fine-tuned in a few-shot setting (Zhou et al., 2022b) with 16 samples per class. The general hyperparameters are listed in Table 7, as per the official implementation. Below, we outline the specific hyperparameters for each VLM tuning method. All methods are adopted from their official implementations. For CoCoOp, no additional hyperparameters are required. In the case of ProDA, we set $\lambda=0.1$. For KgCoOp, $\lambda$ is set to 8.0. MaPLe is configured with a prompt depth $J=0$, and both the language and vision prompt lengths are set to 2. For ProGrad, $\lambda=0.8$ is used. For PromptSRC, deep prompting is set with $V=T=4$, and weight loss employs $\lambda_{1}=10$ and $\lambda_{2}=25$. Finally, for textual diversity, we use $N=60$ standard prompt templates. All methods are reproduced using the standard codebase.
@@ Appendix B The Impact of Different backbones
Following DAC (Wang et al., 2024) experimental settings, we evaluated the performance of DAC and CAC on fine-tuned VLMs with different architectures using the Oxford Flowers102 (Nilsback & Zisserman, 2008) dataset in this section. In contrast to accuracy, which typically exhibits an upward trend with the increase in the number of shots as reported in  (Zhou et al., 2022a), calibration errors are widespread across a variety of few-shot settings. As shown in Table 8, CAC demonstrates stronger cross-model generalization capabilities, achieving effective calibration across all architectures.
Table 8: Comparison results of ECE (%) using different visual backbones on Oxford Flowers102 dataset. The smaller values are better.
@@ Appendix C Detailed Experimental Results
In this section, we provide a comprehensive details of the experimental results, with a focus on the calibration and performance of the fine-tuned VLMs. First, we evaluate the calibration performance on both train and unseen classes, comparing the fine-tuning results. Next, we further explore the calibration performance, discussing how advanced calibration techniques enhance the model’s reliability in making accurate predictions. Additionally, we delve into the relationship between accuracy and confidence. Finally, we present detailed results from the main experiments, showcasing the model’s performance across various tasks and highlighting the improvements achieved through CAC.
@@ C.1 The calibration performance of fine-tuned CLIP
To illustrate the miscalibration in fine-tuned VLMs, we fine-tune the pre-trained CLIP using 5 different tuning methods across 11 downstream datasets, evaluating calibration performance with ECE. In this section, the detailed calibration results are provided. As can be seen from Table 9(b), it is evident that distillation-based methods (KgCoOp, ProGrad, and PromptSRC) play a role in restricting the model’s confidence regarding train classes. For unseen classes, distillation-based methods leverage zero-shot CLIP as a teacher model to enhance generalization, leading to more reliable predictions for these unseen classes, as expected. Through comparison, we find that fine-tuned VLMs tend to be underconfident on the train classes, while they are often overconfident on unseen classes. This observation motivates a deeper investigation into the calibration of fine-tuned VLMs.
Table 9: Expected Calibration Error (ECE) on diverse downstream datasets using various tuning methods for CLIP-ViT-B/16. The calibration performance is averaged across three variants.
(a) Train
@@ C.2 Calibrated performance for fine-tuned CLIP
In our study, we apply commonly employed post-hoc confidence calibration techniques to fine-tuned VLMs, aiming to calibrate both train and unseen classes. Our findings reveal that the established calibration methods are capable of rectifying the miscalibration issues within the train classes. Nevertheless, this effectiveness fails to extend to the unseen classes.
@@ C.2.1 TRAIN CLASS
Post-hoc calibration can remedy miscalibration in train classes. Recent studies (Zadrozny & Elkan, 2001, 2002) have successfully applied scaling-based methods like Temperature Scaling (TS) for VLM calibration, showing that such methods effectively address miscalibration in close-world environments. In addition, we observe that bin-based calibration methods, such as MIR, significantly reduce miscalibration in train classes. CAC, a calibration method designed specifically for VLMs, also generalizes well to train classes and achieves optimal calibration.
Post-hoc calibration for train classes does not extend to unseen classes. For instance, scaling-based methods like TS adjust logits by a single scalar temperature value. When a fine-tuned VLM exhibits underconfidence in the train classes, TS sharpens the logits before the softmax function, increasing confidence in predictions. However, this comes at the cost of making the confidence distribution more peaked, which leads to overconfidence in unseen classes. Bin-based methods, on the other hand, require probabilities from train classes as input, which is incompatible with zero-probability predictions. This validates our motivation and highlights the rationale behind leveraging zero-shot CLIP in our approach.
@@ C.2.2 UNSEEN CLASS
Table 10: Expected Calibration Error (ECE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViT-B/16, with calibration performance averaged across three variants.
@@
Table 11: Expected Calibration Error (ECE) on train classes in diverse downstream datasets using various tuning methods for CLIP-ViT-B/16, with calibration performance averaged across three variants.
@@ C.3 Accuracy & Confidence of fine-tuned CLIP
This section is dedicated to the analysis of miscalibration in the context of classification performance. Since open-vocabulary classes are not encountered during tuning, some might hypothesize that miscalibration stems from a sharp accuracy drop, leading to a higher ECE. We report the average performance across 11 datasets in Table 12(b). While classification performance on unseen classes may be comparable or lower, fine-tuned VLMs still have higher average predictive confidence than zero-shot CLIP, resulting in a larger ECE. This aligns with CAC results: although fine-tuning boosts accuracy, it also increases ECE.
@@ C.4 Detailed results of main experiment
This section showcases detailed results of calibration for unseen classes, aiming to illustrate that CAC facilitates open-vocabulary calibration in current prompt tuning. For a comprehensive evaluation, we use four standard metrics to assess open-vocabulary confidence calibration: Expected Calibration Error (ECE) (Guo et al., 2017), Maximum Calibration Error (MCE) (Guo et al., 2017), Adaptive Calibration Error (ACE) (Nixon et al., 2019), and Proximity Informed Expected Calibration Error (PIECE) (Xiong et al., 2023). All calibration errors are reported in $10^{-2}$.
Table 14: Expected Calibration Error (ECE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViT-B/16, with calibration performance averaged across three variants.
@@
Table 15: Adaptive Calibration Error (ACE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViT-B/16, with calibration performance averaged across three variants.
@@
Table 16: Maximum Calibration Error (MCE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViT-B/16, with calibration performance averaged across three variants.
@@
Table 17: Proximity-Informed Expected Calibration Error (PIECE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViT-B/16, with calibration performance averaged across three variants.